OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. 3 – 7x + 2x when x = 3 A. –12 B. –30 C. –24 D. –18

OpenStudy (anonymous):

= 3 - 7(3) + 2(3) = 3 - 21 + 6 = -12 A

OpenStudy (anonymous):

Hold on let me work it out

OpenStudy (anonymous):

I was just gonna put that

OpenStudy (anonymous):

3-7x+2x 3-21+2x -18+2•3 -18+6 -12